OpenStudy (anonymous):

Find the length and width of a rectangle, given that width is one-half of the length and the perimeter of the rectangle is 54 cm.

OpenStudy (anonymous):

well,suppose that the length of that rectangle is \(\large x \) and the width is \(\large y \). we have \(\large y = \frac{ x }{ 2 }\) , you know perimeter is \(\large x+x+y+y = 2(x+y) \)

OpenStudy (anonymous):



OpenStudy (anonymous):

and perimeter is 54cm so \( 2(x+y) =53 \)

OpenStudy (anonymous):

can you solve this equation and find x and y ?

OpenStudy (anonymous):

@theviolentredhead

OpenStudy (anonymous):

length is 18 and width is 9?

OpenStudy (anonymous):

yes!